# Bipartite Fluctuations of Critical Fermi Surfaces

### v2

@@ Appendix B Coulomb Gauge and Response Theory
where $\Pi_{L}$ and $\Pi_{T}$ represent the longitudinal and transverse components respectively, while $\mathsf{H}$ governs the Hall response. It is sometimes beneficial to decompose the spatial components of the background/gauge field into longitudinal and transverse components, expressed as $\bm{A}=\bm{A}_{L}+\bm{A}_{T}$ where
@@
It is useful to introduce a scalar $A_{T}$ to represent the transverse fields $A_{i}^{T}$, which is defined by
$A_{i}^{T}(k)=\frac{\varepsilon_{ij}k_{j}}{|\bm{k}|}A_{T}(k),\qquad A_{T}(k)=\frac{\varepsilon_{ij}k_{j}}{|\bm{k}|}A_{i}^{T}(k).$ (67)
Under the Coulomb gauge $\bm{A}^{L}=0$, the response theory can expressed in the basis $(A_{\tau},A_{T})$
$\mathcal{S}[A]=\int\frac{\textrm{d}^{3}k}{(2\pi)^{3}}\frac{-1}{2}(A_{\tau}(-k)\;A_{T}(-k))\left(\begin{array}[]{cc}\Pi^{\tau\tau}(k)&\mathsf{H}(k)|\bm{k}|\\ \mathsf{H}(k)|\bm{k}|&\Pi^{TT}(k)\end{array}\right)\left(\begin{array}[]{c}A_{\tau}(k)\\ A_{T}(k)\end{array}\right),$ (72)
where $\Pi^{\tau\tau}(k)=\frac{|\bm{k}|^{2}}{\omega^{2}}\Pi_{L}(\omega,|\bm{k}|)$ and $\Pi^{TT}(k)=-\Pi_{T}(k)$. Here, $k=(\omega,\bm{k})$ is a collective notion of frequency and momentum. Notice that the Chern-Simons term is
$\int\frac{-\mathtt{i}}{4\pi}A\wedge\textrm{d}A=\int\frac{\textrm{d}^{3}k}{(2\pi)^{3}}\frac{-1}{2}(A_{\tau}(-k)\;A_{T}(-k))\frac{|\bm{k}|}{2\pi}\left(\begin{array}[]{cc}0&1\\ 1&0\end{array}\right)\left(\begin{array}[]{c}A_{\tau}(k)\\ A_{T}(k)\end{array}\right).$ (77)
@@ Appendix C Ioffe-Larkin Composition Rule
We investigate the response function of electrons using the parton theory described in Eq. (18). In the standard RPA approach, integrating out both fermionic and bosonic partons $f$ and $b$ leads to
$\mathcal{S}=\int_{k}\frac{-\Pi_{f}^{\mu\nu}(k)}{2}(a_{\mu}(-k)+e_{f}A_{\mu}(-k))(a_{\nu}(k)+e_{f}A_{\nu}(k))+\frac{-\Pi_{b}^{\mu\nu}(k)}{2}(-a_{\mu}(-k)+e_{b}A_{\mu}(-k))(-a_{\nu}(k)+e_{b}A_{\nu}(k)).$ (78)
To enforce gauge invariance, we integrate out the dynamical gauge fields $a_{\mu}$ at the RPA level. This yields the total response theory $\mathcal{L}=\int_{k}\frac{-1}{2}\Pi^{\mu\nu}(k)A_{\mu}(-k)A_{\nu}(k)$ with
@@
where we have used $e_{f}+e_{b}=1$. Therefore, we find the Ioffe-Larkin composition rule
@@
which is independent of the assignment $(e_{f},e_{b})$ of the global $\textrm{U}(1)$ charge.
Under the Coulomb gauge, the parton response functions $\Pi_{f}$ and $\Pi_{b}$ can be written as
@@
The Ioffe-Larkin rule Eq. (80) leads to the gauge-invariant response
@@
For time-reversal invariant systems such that $\mathsf{H}_{f}=0$ and $\mathsf{H}_{b}=0$, one simply has
@@ Appendix D Scaling of Bipartite Fluctuations
In this appendix, we present the technical details of gauge-invariant regularization schemes used to handle the UV divergence in evaluating Eq. (25). Our starting point is the static structure factor of electrons. At long wavelengths, we assume rotational invariance and consider a generic power-law correlation Eq. (2) in real space. To keep our discussion as general as possible, we allow the value of $\alpha$ to be continuously tuned with $\alpha>2$. It is known that the overall coefficient $C_{0}$ is universal in electronic systems described by CFTs where $\alpha=4$, while it is sensitive to interactions in Landau FLs where $\alpha=3$. In App. D.1, we propose a real-space method that is suitable for both dimensional regularization and UV cut-off schemes.
@@
$\Pi^{\tau\tau}(\tau\rightarrow 0,\bm{k})=\int\textrm{d}^{d}x\frac{C_{0}}{|\bm{x}|^{\alpha}}e^{\mathtt{i}\bm{k}\cdot\bm{x}}=\frac{2^{d-\alpha}\pi^{\frac{d}{2}}\Gamma(\frac{d-\alpha}{2})}{\Gamma(\frac{\alpha}{2})}\frac{C_{0}}{\left|\bm{k}\right|^{d-\alpha}}=\begin{cases}+2\pi C_{0}\log(1/|\bm{k}|)&\alpha=2\\ -2\pi C_{0}|\bm{k}|&\alpha=3\\ -\frac{\pi}{2}C_{0}|\bm{k}|^{2}\log(1/|\bm{k}|)&\alpha=4\\ +\frac{2\pi}{9}C_{0}|\bm{k}|^{3}&\alpha=5\\ +\frac{\pi}{32}C_{0}|\bm{k}|^{4}\log(1/|\bm{k}|)&\alpha=6\\ \qquad\vdots&\quad\vdots\end{cases}$ (84)
where $d=2$. In the case of gapless modes in CFLs, there is a contribution of the form $|\bm{k}|^{3}\log(1/|\bm{k}|)$, which results in an additional term $|\bm{x}|^{-5}\log|\bm{x}|$ alongside Eq. (2) with $\alpha=5$. However, this logarithmic correction does not change the scaling of bipartite fluctuations. This issue is further discussed in App. D.2.
@@ D.1 Real-Space Method
We embed the equal-time density response Eq. (2) into a space-time current-current correlation
@@
such that $\Pi^{\tau\tau}(\tau\rightarrow 0,\bm{x})=\langle\check{J}^{\tau}(\tau,\bm{x})\check{J}^{\tau}(\tau,0)\rangle$. We once again introduce a dual gauge field $\check{a}_{\mu}$ to represent the current $\check{J}^{\mu}=\frac{\mathtt{i}}{2\pi}\varepsilon^{\mu\nu\rho}\partial_{\nu}\check{a}_{\rho}$. The gauge-field propagator $\check{D}_{\mu\nu}(x-y)=\langle\check{a}_{\mu}(x)\check{a}_{\nu}(y)\rangle$ can be written as
$\check{D}_{\mu\nu}^{(\alpha)}(x)$ $=\frac{(2\pi)^{2}C_{0}}{(\alpha-2)^{2}}\frac{1}{|x|^{\alpha-2}}\left((1+\zeta)\delta^{\mu\nu}-\zeta(\alpha-2)\frac{x^{\mu}x^{\nu}}{|x|^{2}}\right)$ (86)
@@
where $|\mathcal{C}|=4L$ is the perimeter of the square. By setting $\alpha=3$ and following the calculations in Eq. (9)-(12) from Ref. [18], we obtain the result for Fermi liquids
@@
The general expression Eq. (91) is informative in understanding the contribution of gapless modes in different phases of matter. In non-local systems where the instantaneous charge correlation decays even slower than Landau Fermi liquids (i.e., $\alpha<3$), the first boundary-law term in Eq. (91) vanishes as $\epsilon$ approaches zero, leaving only the second term, which is independent of $\epsilon$ and scales as $|\mathcal{C}|^{4-\alpha}$. Conversely, when $\alpha>3$, the leading-order term is always given by the boundary law $|\mathcal{C}|\epsilon^{3-\alpha}$, which contains a power-law UV divergence. Another notable observation is that when the instantaneous charge correlation is weaker than in CFTs (i.e., $\alpha>4$), the universal subheading term $|\mathcal{C}|^{4-\alpha}$ vanishes in the large-$|\mathcal{C}|$ limit. As we will see, this holds true for CFLs. In Tab. 1, we summarize our findings under the cut-off regularization scheme.
@@ D.2 Momentum-Space Method
by defining the linear functional over the closed loop $\mathcal{C}=\partial\mathcal{A}$ as
@@
$\int\frac{\textrm{d}^{d}\bm{k}}{(2\pi)^{d}}e^{-\mathtt{i}\bm{k}\cdot\bm{x}}\left|\bm{k}\right|^{\beta}=\frac{2^{\beta}\pi^{\frac{1-d}{2}}\Gamma(\frac{d+\beta}{2})}{\Gamma(\frac{d-1}{2})\Gamma(1-\frac{d+\beta}{2})}\frac{1}{|\bm{x}|^{d+\beta}}=\begin{cases}0&\beta=0\\ -\frac{1}{2\pi}|\bm{x}|^{-3}&\beta=1\\ 0&\beta=2\\ +\frac{9}{2\pi}|\bm{x}|^{-5}&\beta=3\\ 0&\beta=4\\ \qquad\vdots&\quad\vdots\end{cases}$ (98)
@@
=$=\begin{cases}-\frac{1}{2\pi|\bm{x}|^{2}}&\beta=0\\ +\frac{\log|\bm{x}|+\log 2+\gamma_{E}-2}{2\pi|\bm{x}|^{3}}&\beta=1\\ +\frac{2}{\pi|\bm{x}|^{4}}&\beta=2\\ -\frac{3(3\log|\bm{x}|+\log 8+3\gamma_{E}-8)}{2\pi|\bm{x}|^{5}}&\beta=3\\ -\frac{32}{\pi|\bm{x}|^{6}}&\beta=4\\ \qquad\vdots&\quad\vdots\end{cases}$ (99)
where again $d=2$, $\Psi^{(0)}(z)=\frac{\textrm{d}}{\textrm{d}z}\log\Gamma(z)$ denotes the the polygamma function, and $\gamma_{E}$ is the Euler–Mascheroni constant. The cases from Eq. (98) with odd $\beta$ and those from Eq. (99) with even $\beta$ together account for all the cases considered earlier in Eq. (84). The new cases arise in Eq. (99) with odd $\beta$, where the power-law correlation receives a logarithmic correction.
@@
$\begin{cases}|\bm{k}|\quad\Rightarrow\quad 2R\log R&\textrm{Landau FLs and superfluids}\\ |\bm{k}|^{3}\log(1/|\bm{k}|)\quad\Rightarrow\quad\frac{3}{8}\frac{(\log R)^{2}}{R}\rightarrow 0&\textrm{gapless modes in CFLs}\\ |\bm{k}|\log(1/|\bm{k}|)\quad\Rightarrow\quad R(\log R)^{2}&\textrm{certain non-Fermi liquids}\end{cases}.$ (103)
@@
Appendix E Dual Theories for CFL-FL Transition

### v3

@@ Appendix B Response Theory and Coulomb Gauge
where $\Pi_{L}$ and $\Pi_{T}$ describe the longitudinal and transverse components respectively, while $\mathsf{H}$ characterizes the Hall response. Here, $k=(\omega,\bm{k})$ collectively denotes the Matsubara frequency and momentum. The $\mathrm{U}(1)$ Ward identity $k_{\mu}\Pi^{\mu\nu}(k)=0$ is automatically satisfied.
To simplify the analysis, it is useful to decompose the spatial components of the background/gauge field into longitudinal and transverse components, expressed as $\bm{A}=\bm{A}_{L}+\bm{A}_{T}$ where
@@
It is convenient to introduce a scalar field $A_{T}$ to represent the transverse component $A_{i}^{T}$, defined by
$A_{i}^{T}(k)=\frac{\varepsilon_{ij}k_{j}}{|\bm{k}|}A_{T}(k)\quad\textrm{or}\quad A_{T}(k)=\frac{\varepsilon_{ij}k_{j}}{|\bm{k}|}A_{i}^{T}(k).$ (75)
In the Coulomb gauge, where $\bm{A}^{L}=0$, the response theory takes a simpler form when expressed in the $(A_{\tau},A_{T})$ basis
$\mathcal{S}[A]=\int\frac{\textrm{d}^{3}k}{(2\pi)^{3}}\frac{1}{2}(A_{\tau}(-k)\;A_{T}(-k))\left(\begin{array}[]{cc}\Pi^{\tau\tau}(k)&\mathsf{H}(k)|\bm{k}|\\ \mathsf{H}(k)|\bm{k}|&\Pi^{TT}(k)\end{array}\right)\left(\begin{array}[]{c}A_{\tau}(k)\\ A_{T}(k)\end{array}\right)$ (80)
with $\Pi^{\tau\tau}(k)=\frac{|\bm{k}|^{2}}{\omega^{2}}\Pi_{L}(\omega,|\bm{k}|)$ and $\Pi^{TT}(k)=-\Pi_{T}(\omega,|\bm{k}|)$. Notice that the Chern-Simons term is
$\int\frac{-\mathtt{i}}{4\pi}A\wedge\textrm{d}A=\int\frac{\textrm{d}^{3}k}{(2\pi)^{3}}\frac{1}{2}(A_{\tau}(-k)\;A_{T}(-k))\frac{-|\bm{k}|}{2\pi}\left(\begin{array}[]{cc}0&1\\ 1&0\end{array}\right)\left(\begin{array}[]{c}A_{\tau}(k)\\ A_{T}(k)\end{array}\right).$ (85)
@@ Appendix C Ioffe-Larkin Composition Rule
We investigate the response function of electrons using the parton theory described in Eq. (18). In the standard RPA approach, integrating out both fermionic and bosonic partons $f$ and $b$ leads to the effective action
$\mathcal{S}=\int_{k}\frac{\Pi_{f}^{\mu\nu}(k)}{2}(a_{\mu}(-k)+e_{f}A_{\mu}(-k))(a_{\nu}(k)+e_{f}A_{\nu}(k))+\frac{\Pi_{b}^{\mu\nu}(k)}{2}(-a_{\mu}(-k)+e_{b}A_{\mu}(-k))(-a_{\nu}(k)+e_{b}A_{\nu}(k)).$ (86)
To enforce gauge invariance, we integrate out the dynamical gauge fields $a_{\mu}$ at the RPA level. This yields the total response theory $\mathcal{L}=\int_{k}\frac{1}{2}\Pi^{\mu\nu}(k)A_{\mu}(-k)A_{\nu}(k)$ with
@@
where we have used $e_{f}+e_{b}=1$. Therefore, we find the celebrated Ioffe-Larkin composition rule [89]
@@
which is independent of the specific assignment $(e_{f},e_{b})$ of the global $\textrm{U}(1)$ charge among partons.
In the Coulomb gauge, the parton response functions $\Pi_{f}$ and $\Pi_{b}$ can be written as
@@
Applying the Ioffe-Larkin rule Eq. (88), the resulting gauge-invariant response functions are
@@
In time-reversal invariant systems, where $\mathsf{H}_{f}=0$ and $\mathsf{H}_{b}=0$, these simplify further to
@@ Appendix D Scaling of Bipartite Fluctuations
In this appendix, we present the technical details of gauge-invariant regularization schemes used to handle the UV divergence in evaluating Eq. (25). Our starting point is the static structure factor of electrons. At long wavelengths, we assume rotational invariance and consider a generic power-law correlation Eq. (2) in real space. To keep our discussion as general as possible, we allow the value of $\alpha$ to be continuously tuned with $\alpha>2$.
@@
$\int\textrm{d}^{d}x\frac{-C_{0}}{|\bm{x}|^{\alpha}}e^{\mathtt{i}\bm{k}\cdot\bm{x}}=\frac{2^{d-\alpha}\pi^{\frac{d}{2}}\Gamma(\frac{d-\alpha}{2})}{\Gamma(\frac{\alpha}{2})}\frac{-C_{0}}{\left|\bm{k}\right|^{d-\alpha}}=\begin{cases}-2\pi C_{0}\log(1/|\bm{k}|)&\alpha=2\\ +2\pi C_{0}|\bm{k}|&\alpha=3\\ +\frac{\pi}{2}C_{0}|\bm{k}|^{2}\log(1/|\bm{k}|)&\alpha=4\\ -\frac{2\pi}{9}C_{0}|\bm{k}|^{3}&\alpha=5\\ -\frac{\pi}{32}C_{0}|\bm{k}|^{4}\log(1/|\bm{k}|)&\alpha=6\\$ $\qquad\vdots$$\quad\vdots\end{cases}$ (92)
where $d=2$. Although Eq. (92) includes the Fourier relation for the case $\alpha=2$, physical systems with a global U(1) symmetry must satisfy $\alpha>2$, since the static structure is required to vanish in the limit $\bm{k}\rightarrow 0$. This constraint can be understood by noting that, in this limit, the density operator $J^{\tau}(\tau,\bm{k}\rightarrow 0)$ corresponds to the globally conserved charge $Q=\mathtt{i}\int\textrm{d}^{2}\bm{x}J^{\tau}(\tau,\bm{x})$. As a result, the equal-time density-density correlation is simply the charge variance
$\textrm{Var}(Q)=\langle Q^{2}\rangle-\langle Q\rangle^{2}$ (93)
evaluated in the zero-temperature ground state. Since $[Q,H]=0$, the ground state of the Hamiltonian $H$ must also be an eigenstate of $Q$, implying that $\textrm{Var}(Q)=0$.
In the case of gapless modes in CFLs, the static structure factor contains a contribution of the form $|\bm{k}|^{3}\log(1/|\bm{k}|)$, which leads to an additional term $|\bm{x}|^{-5}\log|\bm{x}|$ in real space, alongside Eq. (2) with $\alpha=5$. However, this logarithmic correction does not change the scaling of bipartite fluctuations. This issue is further discussed in App. D.2.
@@ D.1 Real-Space Method
In App. D.1, we propose a real-space method that is suitable for both dimensional regularization and UV cut-off schemes. The idea is to embed the equal-time correlation function Eq. (2) into a space-time current-current correlation
@@
such that $\Pi^{\tau\tau}(\tau\rightarrow 0,\bm{x})=-\langle\check{J}^{\tau}(\tau,\bm{x})\check{J}^{\tau}(\tau,0)\rangle$. We once again introduce a dual gauge field $\check{a}_{\mu}$ to represent the current $\check{J}^{\mu}=\frac{\mathtt{i}}{2\pi}\varepsilon^{\mu\nu\rho}\partial_{\nu}\check{a}_{\rho}$. The gauge-field propagator $\check{D}_{\mu\nu}(x-y)=\langle\check{a}_{\mu}(x)\check{a}_{\nu}(y)\rangle$ can be written as
$\check{D}_{\mu\nu}^{(\alpha)}(x)$ $=\frac{(2\pi)^{2}C_{0}}{(\alpha-2)^{2}}\frac{1}{|x|^{\alpha-2}}\left((1+\zeta)\delta^{\mu\nu}-\zeta(\alpha-2)\frac{x^{\mu}x^{\nu}}{|x|^{2}}\right)$
@@
where $|\mathcal{C}|=4L$ is the perimeter of the square. By setting $\alpha=3$ and following the calculations in Eq. (9)-(12) from Ref. [18], we obtain the result for Landau FLs
@@
The general expression Eq. (99) is informative in understanding the contribution of gapless modes in different phases of matter. In non-local systems where the instantaneous charge correlation decays even slower than Landau FLs (i.e., $2<\alpha<3$), the first boundary-law term in Eq. (99) vanishes as $\epsilon$ approaches zero, leaving only the second term, which is independent of $\epsilon$ and scales as $|\mathcal{C}|^{4-\alpha}$. Conversely, when $\alpha>3$, the leading-order term is always given by the boundary law $|\mathcal{C}|\epsilon^{3-\alpha}$, which contains a power-law UV divergence. Another notable observation is that when the instantaneous charge correlation is weaker than in CFTs (i.e., $\alpha>4$), the universal subheading term $|\mathcal{C}|^{4-\alpha}$ vanishes in the large-$|\mathcal{C}|$ limit. As we will see, this holds true for CFLs. In Tab. 1, we summarize our findings under the cut-off regularization scheme.
@@ D.2 Momentum-Space Method
where $Y_{\mathcal{C}}^{i}(\bm{k})$ is a linear functional over the closed loop $\mathcal{C}=\partial\mathcal{A}$, defined by
@@
$\int\frac{\textrm{d}^{d}\bm{k}}{(2\pi)^{d}}e^{-\mathtt{i}\bm{k}\cdot\bm{x}}\left|\bm{k}\right|^{\beta}=\frac{2^{\beta}\pi^{\frac{1-d}{2}}\Gamma(\frac{d+\beta}{2})}{\Gamma(\frac{d-1}{2})\Gamma(1-\frac{d+\beta}{2})}\frac{1}{|\bm{x}|^{d+\beta}}=\begin{cases}0&\beta=0\\ -\frac{1}{2\pi}|\bm{x}|^{-3}&\beta=1\\ 0&\beta=2\\ +\frac{9}{2\pi}|\bm{x}|^{-5}&\beta=3\\ 0&\beta=4\\$ $\qquad\vdots$$\quad\vdots\end{cases}$ (106)
@@
$=\begin{cases}-\frac{1}{2\pi|\bm{x}|^{2}}&\beta=0\\ +\frac{\log|\bm{x}|+\log 2+\gamma_{E}-2}{2\pi|\bm{x}|^{3}}&\beta=1\\ +\frac{2}{\pi|\bm{x}|^{4}}&\beta=2\\ -\frac{3(3\log|\bm{x}|+\log 8+3\gamma_{E}-8)}{2\pi|\bm{x}|^{5}}&\beta=3\\ -\frac{32}{\pi|\bm{x}|^{6}}&\beta=4\\$ $\qquad\vdots$$\quad\vdots\end{cases}$ (107)
where again $d=2$, $\Psi^{(0)}(z)=\frac{\textrm{d}}{\textrm{d}z}\log\Gamma(z)$ denotes the the polygamma function, and $\gamma_{E}$ is the Euler-Mascheroni constant. The cases from Eq. (106) with odd $\beta$ and those from Eq. (107) with even $\beta$ together account for all the cases considered earlier in Eq. (92). The new cases arise in Eq. (107) with odd $\beta$, where the power-law correlation receives a logarithmic correction.
@@
[…]
@@
Appendix E Optical Absorption and Structure Factor
For gapped charge insulators, the topological lower bound in Eq. (55) can be understood from the positive semi-definiteness of the many-body quantum geometric tensor, where the quantum metric and Berry curvature correspond to its real and imaginary parts, respectively [64, 65, 66, 71, 83]. For gapless charge insulators (with a finite $\mathsf{G}^{ij}$), such as CFLs, we provide an additional argument based on the positivity of optical absorption power, directly following the discussion in Ref. [82, 84] for gapped states.
We perturb the system with a monochromatic electric field $\bm{E}(t)=\mathcal{E}^{*}e^{\mathtt{i}\omega t}+\mathcal{E}e^{-\mathtt{i}\omega t}$. The absorption power is
$P\sim\int\textrm{d}t\bm{j}(t)\cdot\bm{E}(t)=\mathcal{E}^{*}\cdot\bm{j}(\omega)+\mathcal{E}\cdot\bm{j}^{*}(\omega)=\mathcal{E}_{i}^{*}(\sigma^{ij}(\omega)+\sigma^{*ji}(\omega))\mathcal{E}_{j}.$ (112)
It is convenient to define the absorptive part of the conductivity as
$\sigma_{\textrm{abs}}^{ij}(\omega)=\frac{\sigma^{ij}(\omega)+\sigma^{*ji}(\omega)}{2}=\textrm{Re}\sigma_{+}^{ij}(\omega)+\mathtt{i}\textrm{Im}\sigma_{-}^{ij}(\omega),$ (113)
and the associated optical weight as
$W_{[\ell]}^{ij}=\int_{0}^{+\infty}\frac{\textrm{d}\omega}{\pi}\frac{\sigma_{\textrm{abs}}^{ij}(\omega)}{\omega^{\ell}}.$ (114)
Here, we use $\sigma_{\pm}^{ij}=(\sigma^{ij}\pm\sigma^{ji})/2$ to denote the longitudinal and Hall conductivities, respectively. Assuming the unperturbed system is in the ground state, the absorption power given by Eq. 112 is expected to be non-negative, i.e., $P\sim\mathcal{E}_{i}^{*}\sigma_{\textrm{abs}}^{ij}(\omega)\mathcal{E}_{j}\geq 0$ holds for any complex vector $\mathcal{E}$. Consequently, both $\sigma_{\textrm{abs}}^{ij}(\omega)$ at each frequency and the optical weight $W_{[\ell]}^{ij}$, as given in Eq. 114, are positive semi-definite matrices.
The real part of $W_{[1]}^{ij}$ is given by the SWM sum rule in Eq. (10), while the imaginary part of $W_{[1]}^{ij}$ is related to the DC Hall conductivity $\sigma^{xy}=\textrm{Re}\sigma_{-}^{xy}(0)$ through the Kramers-Kronig relation
$\textrm{Re}\sigma_{-}^{ij}(\omega)=\frac{2}{\pi}\int_{0}^{+\infty}\textrm{d}\nu\frac{\nu}{\nu^{2}-\omega^{2}}\textrm{Im}\sigma_{-}^{ij}(\nu).$ (115)
In conclusion, one obtains the positive semi-definite matrix
$0\leq W_{[1]}^{ij}=\mathsf{G}^{ij}+\frac{\mathtt{i}}{2}\textrm{Re}\sigma_{-}^{ij}(0),$ (116)
which corresponds to the lower bound stated in Eq. 55. In two dimensions, using $\det(W_{[1]})\geq 0$, one can show that
$\frac{\textrm{Tr}(\mathsf{G})}{2}\geq\sqrt{\det(\mathsf{G})}\geq\frac{|\sigma^{xy}|}{2}.$ (117)
For gapped charge insulators, $W_{[1]}^{ab}$ is precisely the many-body quantum geometric tensor, as defined by flux insertion or twisted boundary conditions. Importantly, the lower bound is expected to hold as long as the absorption power in Eq. 112 remains non-negative, even if the system is gapless and the DC Hall conductivity $\sigma_{xy}$ is not a topological invariant. For both gapless and gapped QH states, Ref. [71] clarifies the conditions for bound saturation from two perspectives: (1) invariance under continuous translations and Galilean boosts, and (2) the holomorphic properties of many-body wavefunctions, including Laughlin wavefunctions and composite-fermion Fermi sea wavefunctions.
Appendix F Dual Theories for CFL-FL Transition